OpenStudy (anonymous):

solve the proportion 35 over 28 = x over 4. what is the value of x 3 5 7 11

OpenStudy (anonymous):

\[\Large \frac{35}{28}=\frac{x}{4} \] That's it ?

OpenStudy (anonymous):

Use Cross Multiplication then solve for x .

OpenStudy (anonymous):

For this problem you need to cross multiply. 35*4 = 28x Then divide by 28 for both sides. What do you get as x?

OpenStudy (anonymous):

\[\Large 4 \times 35 = x\times 28 \] \[\Large x=\frac{4 \times 35}{28}\]

OpenStudy (anonymous):

Calculate it and find the answer :)

OpenStudy (anonymous):

Got it ?

OpenStudy (anonymous):

Click Best Response

OpenStudy (anonymous):

I think its 5